b) X-N(-8,12). Find: i P(X<-9.8) ii. P(X > -8.2) iii. P(-7< X <0.5)
Answer #1

Solution :

i

P(x < -9.8) = P[(x - ) / < (-9.8 - (-8)) / ]

= P(z < -0.15)

= 0.4404

P(x < -9.8) = 0.4404

ii

P(x > -8.2) = 1 - P(x < -8.2)

= 1 - P[(x - ) / < (-8.2 - (-8)) / 12)

= 1 - P(z < -0.0167)

= 1 - 0.4933

= 0.5067

P(x > -8.2) = 0.5067

iii

P(-7 < x < 0.5) = P[(-7 - (-8))/ 12) < (x - ) /  < (0.5 - (-8)) / 12) ]

= P(0.0833 < z < 0.7083)

= P(z < 0.7083) - P(z < 0.0833)

= 0.7606 - 0.5332

= 0.2274

P(-7 < x < 0.5) = 0.2274